What is the rule?
You have a table with the following:
n y
2 8
7 3
5 5
What is the linear equation?
y = ?
kspjrp@gmail.com

Answer by stanbon(75887)

--------------------
You have a table with the following:
n y
2 8
5 5
7 3
----
Slope between (2,8) and (5,5) is (5-8)/(5-2) = -1
Slope between (5,5) and (7,3) is (3-5)/(7-5) = -1
What is the linear equation?
Form:
y = mn + b
5 = (-1)5 + b
b = 10
----
y = -n + 10
==========================
Cheers,
Stan H.
==============